# Determinant Estimation under Memory Constraints and Neural Scaling Laws

### v1

@@ Appendix F Background Material for Neural Scaling Laws
where $\mu$ is the marginal distribution on $\mathcal{X}$ of $\rho$, the source distribution on $(\mathcal{X}\times\mathcal{Y})$ underlying the dataset. The operator $T$ can be decomposed by the spectral theorem of compact self-adjoint operators via
@@
$[\mathcal{H}]^{s}=\mathrm{Range}(T^{s/2})=\left\{\sum_{i=1}^{\infty}a_{i}\lambda_{i}^{\frac{s}{2}}e_{i}\mid\sum_{i=1}^{\infty}a_{i}^{2}<\infty\right\}\subset L^{2}.$ (F.5)
@@
Finally, we have the following assumption on the smoothness of the source function $f^{\star}_{\rho}=\mathbb{E}_{\rho}[y\mid x]$, which is a more precise characterization than requiring it to belong to some interpolation space.
@@
Appendix G Proofs of Lemma 1 and Proposition 1
@@
    Fix $\alpha>0$ and write $\mathbf{K}_{n}^{\alpha}=\mathbf{K}_{n}+\alpha\mathbf{Id}_{n\times n}$. In block form, $\mathbf{K}_{n}^{\alpha}$ contains $\mathbf{K}_{n-1}^{\alpha}$ according to
@@
    where $\kappa(\boldsymbol{x}_{n-1},x_{n})=(\kappa(x_{i},x_{n}))_{i=1}^{n-1}\in\mathbb{R}^{nd\times d}$. Consequently, since $\mathbf{K}_{n}^{\alpha}$ and $\mathbf{K}_{n-1}^{\alpha}$ are both positive-definite, their determinants differ by the Schur determinant:
@@
    $\mathsf{pdet}(\mathbf{K}_{n})\leq\mathsf{pdet}(\mathbf{K}_{n-1})\;\mathsf{pdet}(\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)).$ (G.1)
@@
    Letting $c_{0}\coloneqq\log C-\mathsf{logdet}(\mathbf{K}_{1})$ and dividing by $n$ implies the first result. For the second result, we replace the $o_{p}(1)$ term in (G.2) with $\delta_{n-1}$. Consequently,
@@
Appendix H Derivation of FLODANCE Parameterization
Letting $\nu_{n}=\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$, Proposition 1 implies that, asymptotically in $n$,
@@ Appendix H Derivation of FLODANCE Parameterization
and equivalently,
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n},$
where
$x_{n,0}=\sqrt{n-1},\qquad x_{n,i}=\frac{\log(n!)}{n^{i-1}\sqrt{n-1}},\qquad y_{n}=\frac{n}{\sqrt{n-1}}(L_{n}-L_{1}).$
@@
To validate the theoretical parameterization, we empirically compute the log-determinants of NTK submatrices of size $1,\dots,n$ from ResNet9 and ResNet50 trained on CIFAR-10, where $n=$50000$$. The log-determinants are computed using MEMDET (Algorithm 3) with LDL decomposition. Figure H.1 presents the empirical results, where the black curves (overlaid by orange curves) represent the computed log-determinants, and the orange curves show the theoretical fits obtained using the parameterization in Algorithm 1. The fitting is performed globally over the entire interval, demonstrating the effectiveness of the theoretical model in capturing the log-determinant behavior. The near-perfect overlap between the curves in both cases highlights the accuracy of the fit, with errors remaining below 0.05% for most of the interval.
@@
Figure H.1: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). The orange curves represent theoretical fits based on the parametrization derived in Algorithm 1. Fitting is performed globally over the entire interval $(n_{0},n)=(1,5\times 10^{4})$, demonstrating the accuracy of the theoretical model. The blue curve, corresponding to the right axis (scaled to one-thousandth of the left axis), shows the absolute error.
@@ Appendix I Implementation and Reproducibility Guide
We developed a Python package detkit444detkit is available for installation from PyPI (https://pypi.org/project/detkit), the documentation can be found at https://ameli.github.io/detkit, and the source code is available at https://github.com/ameli/detkit. that implements the MEMDET algorithm and can be used to reproduce the numerical results of this paper. A minimalistic usage of the detkit.memdet function is shown in LABEL:list:detkit1, where the user can specify various parameters: the maximum memory limit (max_mem), the structure of the matrix via the assume argument—set to gen for generic matrices (Algorithm 2), sym for symmetric matrices (Algorithm 3), and spd for symmetric positive-definite matrices (Algorithm 4)—whether the data is provided in full or in its lower/upper triangular form (triangle), the arithmetic precision used during computation (mixed_precision), the location of scratchpad space on disk (scratch_dir), and enabling parallel data transfer between memory and disk (parallel_io).
@@
6M = zarr.open(’filename.zarr’, ’r’)
@@
An example of utilizing FLODANCE is shown in LABEL:list:detkit2 using the FitLogdet class in detkit. A small subset of log-determinants from submatrices, computed in LABEL:list:detkit1, is used to extrapolate log-determinants of larger submatrices via Algorithm 1.
@@
We have also concurrently developed a separate high-performance Python package, imate,555imate is available for installation from PyPI (https://pypi.org/project/imate), the documentation can be found at https://ameli.github.io/imate, and the source code is available at https://github.com/ameli/imate. which implements stochastic Lanczos quadrature (SLQ), a randomized method for approximating the log-determinant at scale. This package is implemented with a C++/CUDA backend and supports execution on both CPU and multiple GPUs. LABEL:list:imate demonstrates a minimalistic usage of the imate.logdet function.
@@
# Number of detests (;#\$ n \$#;) and labels (;#\$ d \$#;)

### v2

@@ F.1 Background Material for Neural Scaling Laws
where $\mu$ is the marginal distribution of $\rho$ on $\mathcal{X}$, where $\rho$ is the source distribution on $(\mathcal{X}\times\mathcal{Y})$ underlying the dataset. The operator $T$ can be decomposed by the spectral theorem of compact self-adjoint operators via
@@
$[\mathcal{H}]^{s}=\mathrm{range}(T^{s/2})=\left\{\sum_{i=1}^{\infty}a_{i}\lambda_{i}^{\frac{s}{2}}e_{i}\,\middle|\,\sum_{i=1}^{\infty}a_{i}^{2}<\infty\right\}\subset L^{2}.$ (F.5)
@@
Finally, we have the following assumption on the smoothness of the source function $f^{\star}_{\rho}=\mathbb{E}_{\rho}[y\,|\,x]$, which is a more precise characterization than requiring it to belong to some interpolation space.
@@
F.2 Proofs of Lemmas 1 and 1
@@
    Fix $\alpha>0$ and write $\mathbf{K}_{n}^{\alpha}=\mathbf{K}_{n}+\alpha\mathbf{I}$ where $\mathbf{I}$ is the identity matrix. In block form, $\mathbf{K}_{n}^{\alpha}$ contains $\mathbf{K}_{n-1}^{\alpha}$ according to
@@
    where $\kappa(\boldsymbol{x}_{n-1},x_{n})=(\kappa(x_{i},x_{n}))_{i=1}^{n-1}\in\mathbb{R}^{(n-1)d\times d}$. Consequently, since $\mathbf{K}_{n}^{\alpha}$ and $\mathbf{K}_{n-1}^{\alpha}$ are both positive-definite, their determinants differ by the Schur determinant:
@@
    $\mathsf{pdet}(\mathbf{K}_{n})\leq\mathsf{pdet}(\mathbf{K}_{n-1})\;\mathsf{pdet}(\mathrm{Cov}(f(x_{n})~{}|~{}f(x_{i})=0\text{ for }i=1,\dots,n-1)).$ (F.6)
@@
    Letting $c_{0}\coloneqq\log C-\mathsf{logdet}(\mathbf{K}_{1})$ and dividing by $n$ implies the first result. For the second result, we replace the $o_{p}(1)$ term in (F.7) with $\delta_{n-1}$. Consequently,
@@
F.3 Derivation of FLODANCE Parameterization
Here we derive the equation (6), which leads to the numerical procedure in Algorithm 1. To show that the coefficients $c_{0}$ and $\nu_{0},\dots,\nu_{q}$ can be obtained using standard linear regression procedures, it is necessary to determine the form of the covariates $x_{n,i}$. Letting $\nu_{n}\coloneqq\nu_{0}+\sum_{i=1}^{q}\nu_{i}n^{-i}$, Proposition 1 implies that, asymptotically in $n$,
@@ F.3 Derivation of FLODANCE Parameterization
Equivalently, the above relation can be recast as a linear regression problem:
$y_{n}=c_{0}x_{n,0}+\sum_{i=1}^{q+1}\nu_{i-1}x_{n,i}+\epsilon_{n}.$
where $c_{0},\nu_{0},\dots,\nu_{q}$ are the regression coefficients for the covariates $x_{n,i}$ defined as
$x_{n,i}=\begin{cases}\sqrt{n-1},&i=0,\\ \displaystyle{\frac{\log(n!)}{n^{i-1}\sqrt{n-1}}},&i=1,\dots q.\end{cases}$
The target variable for regression is given by
$y_{n}=\frac{n}{\sqrt{n-1}}(L_{n}-L_{1}).$
We note that in numerical implementations, the term $\log(n!)$ should be evaluated using the log-gamma function: $\log\Gamma(n+1)=\log(n!)$.
Appendix G Further Empirical Analysis of FLODANCE
This section investigates the accuracy, robustness, and generality of FLODANCE through two complementary studies. Section G.1 revisits NTK matrices, quantifying the method’s sensitivity to fitting-interval length and to random subsampling. Section G.2 then applies FLODANCE to a multi-output Gaussian process with a Matérn kernel, demonstrating that the same scaling-law machinery applies well beyond neural kernels.
G.1 Sensitivity and Robustness on NTK Matrices
Global Fit of FLODANCE on NTK Matrices.
@@ G.1 Sensitivity and Robustness on NTK Matrices
To validate the theoretical parameterization given in Section F.3, we compute the log-determinants of NTK submatrices of size $1,\dots,n$ from ResNet9 and ResNet50 trained on CIFAR-10, with $n=$50000$$. The log-determinants are obtained using MEMDET (Algorithm 3) with LDL decomposition. Figure G.1 presents the empirical results: black curves (largely overlaid by orange) show the computed log-determinants, while the orange curves show the theoretical fits derived from the parameterization in Algorithm 1. The fitting is performed globally over the entire interval, demonstrating the accuracy of the theoretical model in capturing the log-determinant behavior. This global fit complements the extrapolation-based application shown earlier in Figure 5, where FLODANCE is trained on a small subset and extrapolated to the full range. The near-perfect overlap between the curves highlights the quality of the fit, with errors remaining below 0.05% for most of the interval.
@@
Figure G.1: Log-determinant $\ell_{n}$ for $n=1,\dots,$50000$$, corresponding to $m\times m$ NTK submatrices where $m=nd$ and $d=10$, from 64-bit NTK matrices of ResNet9 (a) and ResNet50 (b) trained on CIFAR-10 with $50000$ datapoints. Values are computed using MEMDET (Algorithm 3) with LDL decomposition (black curves, overlaid by colored curves). The orange curves represent theoretical fits based on the parametrization derived in Algorithm 1. Fitting is performed globally over the entire interval $(n_{0},n)=(1,5\times 10^{4})$, demonstrating the accuracy of the theoretical model. The blue curve, corresponding to the right axis (scaled to one-thousandth of the left axis), shows the absolute error.
Uncertainty Under Subsampling.
To assess the robustness of FLODANCE predictions, we evaluate how subsampling affects log-determinant estimates. In Figure G.2, we generate 15 independent subsamples of NTK matrices of size $m=nd$, with $d=10$ and $n=$10000$$, from ResNet50 trained on CIFAR-10. Panel (a) shows the exact log-determinants $\ell_{n}$ computed using MEMDET for each subsample. The black curve denotes the ensemble mean, while the gray shading around the mean (barely visible) indicates the standard deviation across subsamples. The red curve (right axis of panel) shows the normalized standard deviation, which remains below $0.1\%$ throughout, indicating remarkable stability of the log-determinants under subsampling.
Panel (b) evaluates the predictive performance of FLODANCE on the same ensembles, where the model is trained on the interval $(n_{0},n_{s})=(1,10^{3})$ and extrapolated to $(n_{s},n)=(10^{3},10^{4})$. The predicted mean log-determinant is shown in red (extrapolated) and yellow (fitted), closely tracking the ensemble mean of the exact log-determinants (black, largely obscured). The right axis displays the relative error: the blue curve denotes the mean, which stays below 0.5%, and the shaded region shows the standard deviation across ensembles, which remains comparably tight. These results confirm that FLODANCE exhibits both accuracy and robustness under random subsampling.
Figure G.2: Sensitivity of log-determinant estimates to subsampling variation. (a) Exact log-determinants $\ell_{n}$ of 15 randomly subsampled NTK matrices of size $m\times m$, $m=nd$, with number of classes $d=10$ and data points $n=$10000$$, computed using MEMDET on ResNet50 trained on CIFAR-10. The black curve denotes the mean, and the shaded gray (barely visible) shows the standard deviation across subsamples. The right ordinate shows the normalized standard deviation (red), which remains below 0.1%. (b) Predicted log-determinants using FLODANCE fitted over a small interval $(n_{0},n_{s})=(1,10^{3})$ (yellow), and extrapolated to $(n_{s},n)=(10^{3},10^{4})$ (red). The left axis shows the predicted mean; the right axis shows the relative error (blue), with mean error under 1% and variation across ensemble (shaded blue).
Effect of Fitting Interval Size.
We investigate the sensitivity of FLODANCE to the choice of the fitting interval size $n_{s}$, which governs the trade-off between computational cost and extrapolation accuracy. Recall that FLODANCE fits a model to log-determinants computed over the interval $[n_{0},n_{s}]$ and extrapolates to the full range $[n_{s},n]$, where $n_{0}=1$ and $n=$50000$$ in this experiment. Figure G.3 evaluates this trade-off on NTK matrices from ResNet50 trained on CIFAR-10. As $n_{s}$ increases, more data is used for fitting (increasing the cost), but less extrapolation is required (increasing the accuracy).
Panel (a) shows the root-mean-square (RMS) error of the fit in the training interval $[n_{0},n_{s}]$, which increases with $n_{s}$ due to the growing number of points to match. Panel (b) displays the RMS error in the extrapolation region $[n_{s},n]$, which decreases as the extrapolation range shrinks. Finally, (c) plots the relative error of the prediction at the endpoint $n=$50000$$, which drops sharply from over $50\%$ to under $0.1\%$ as $n_{s}$ increases. These results illustrate the central design principle of FLODANCE: by choosing a moderate $n_{s}$, one can achieve accurate predictions while keeping the cost of computing exact log-determinants limited to smaller submatrices.
Figure G.3: Sensitivity of FLODANCE to the choice of fitting interval size $n_{s}$. Based on ResNet50 trained on the full CIFAR-10 dataset with $n=5\times 10^{4}$ data points and $d=10$ classes, resulting in NTK matrices of size $m=nd=10^{5}$. FLODANCE extrapolates log-determinants by fitting a model on submatrices of size $n_{0}=1$ to $n_{s}$, and extending this fit to larger sizes up to $n$. (a) Root-mean-square error (RMSE) of the fit in the interval $[1,n_{s}]$, showing increasing fitting error as $n_{s}$ grows. (b) RMSE of extrapolation in the interval $[n_{s},n]$, which decreases with $n_{s}$. (c) Relative error of predicting the log-determinant at $n=5\times 10^{4}$, again decreasing with $n_{s}$.
G.2 Extension to Matérn Kernel Gaussian Process
Finally, we demonstrate the generality of FLODANCE beyond NTKs by applying it to a multi-output Gaussian process with a Matérn kernel, widely used in spatial statistics due to its tunable smoothness. The isotropic Matérn correlation function of Matérn, (1960) (see also Stein, (1999, p. 31)) between two spatial points $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{p}$ is given by
$\rho(\boldsymbol{x},\boldsymbol{x}^{\prime}\,|\,\alpha,\nu)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\sqrt{2\nu}\,\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{2}}{\alpha}\right)^{\nu}K_{\nu}\left(\sqrt{2\nu}\,\frac{\|\boldsymbol{x}-\boldsymbol{x}^{\prime}\|_{2}}{\alpha}\right),$
where $\Gamma(\cdot)$ is the Gamma function and $K_{\nu}(\cdot)$ is the modified Bessel function of the second kind of order $\nu$ (Abramowitz & Stegun, , 1964, Section 9.6). The hyperparameter $\nu$ modulates the smoothness of the underlying random process, and the hyperparameter $\alpha>0$ is the correlation scale of the kernel. We construct the multi-output covariance using the linear model of coregionalization (LMC) (Gelfand et al., , 2010, Section 28.7) given by
$\kappa(\boldsymbol{x},\boldsymbol{x}^{\prime}\,|\,\alpha,\nu)=\sigma(\boldsymbol{x})^{\frac{1}{2}}\sigma(\boldsymbol{x}^{\prime})^{\frac{1}{2}}\rho(\boldsymbol{x},\boldsymbol{x}^{\prime}\,|\,\alpha,\nu),$
where $\sigma:\mathbb{R}^{p}\to\mathbb{R}^{d\times d}$ is the local covariance of the model’s vector output of size $d$. In this model, we use the matrix square root, $\sigma^{\frac{1}{2}}$, to project the scalar Matérn correlation into the $d\times d$ coregionalization space while preserving positive-definiteness.
In our experiment, we generated $n=$10000$$ random spatial points in dimension $p=2$ for a Gaussian process with output dimension $d=10$, resulting in a covariance matrix of size $m=nd=$100000$$. We set the Matérn correlation scale to $\alpha=0.04$ and the smoothness parameter to $\nu=1.5$. The local covariance fields $\sigma(\boldsymbol{x})$ were instantiated as random symmetric positive-definite matrices drawn from a Wishart distribution.
Figure G.4: Application of FLODANCE to a multi-output Gaussian process with a Matérn kernel. We generate $n=$10000$$ spatial locations in $\mathbb{R}^{2}$ and assume a $d=10$-dimensional output per location, resulting in a covariance matrix of size $m=nd=$100000$$. The covariance structure follows a Matérn kernel with smoothness $\nu=1.5$ and scale parameter $\alpha=0.04$, combined with a linear model of coregionalization (LMC) for output covariances. (a) Scale law illustrated by the ratio of successive determinants over increasing submatrix sizes. (b) Log-determinant prediction using FLODANCE. The black curve (left axis, largely obscured) is the exact log-determinant $\ell_{n}$ computed by MEMDET. FLODANCE is fitted on $[1,n_{s}=10^{3}]$ (yellow) and extrapolated to $[n_{s},n=10^{4}]$ (red). The blue curve (right axis) shows the relative error of prediction, which remains below 0.4%.
Figure G.4 illustrates the effectiveness of FLODANCE on this Matérn-based covariance structure. Panel (a) shows the empirical scale law via the ratio of successive log-determinants, which exhibits a smooth trend consistent with the theoretical behavior observed for NTKs. Panel (b) evaluates the extrapolation accuracy of FLODANCE: the black curve denotes the exact log-determinant $\ell_{n}$ computed via MEMDET, while the colored curves show the fit (yellow) over the small interval $[1,n_{s}=10^{3}]$ and the extrapolation (red) to $[n_{s},n=10^{4}]$. The right axis displays the relative error (blue), which remains below $0.4\%$ throughout. This example demonstrates FLODANCE’s flexibility in handling structured kernels beyond neural tangent models.
Appendix H Comparison of Log-Determinant Methods: Complexity and Runtime
We summarize the computational characteristics of the log-determinant approximation methods evaluated in this work. Table H.1 compares their theoretical computational complexity, highlighting how each method accesses or approximates the full matrix—whether through exact computation with full matrix access (MEMDET), or through approximation strategies such as subsampling (FLODANCE), matrix–vector product oracles (SLQ), or blockwise approximations (Pseudo NTK and Block Diagonal).
Table H.1: Computational complexity of log-determinant estimation methods. The first row corresponds to the exact method (MEMDET), while all other methods are approximations. One FLOP is counted as a fused multiply-add (FMA) operation.
Method Approach Complexity Description MEMDET Direct factorization $\frac{1}{6}m^{3}-\frac{1}{4}m^{2}+\frac{1}{12}m$ $m$ : Full matrix size, $m=nd$ FLODANCE Submatrix extrapolation $\frac{1}{6}m_{s}^{3}-\frac{1}{4}m_{s}^{2}+\frac{1}{12}m_{s}+(q+3)^{2}n_{s}$ $n_{s}$ : Number of data samples from $n$ $m_{s}$ : Sampled matrix size, $m_{s}=n_{s}d$ $q$ : Laurent series truncation order SLQ Stochastic trace estimation $(m^{2}l+ml^{2})s$ $l$ : Krylov subspace size $s$ : Number of Monte Carlo samples Pseudo NTK Cross-class block reduction $\frac{1}{6}\left(\frac{m}{d}\right)^{3}-\frac{1}{4}\left(\frac{m}{d}\right)^{2}+\frac{1}{12}\left(\frac{m}{d}\right)+m^{2}$ $m$ : Full matrix size $d$ : Number of model outputs Block Diagonal Class-wise block approx. $\frac{1}{6}md^{2}-\frac{1}{4}md+\frac{1}{12}m$ $m$ : Full matrix size $d$ : Number of model outputs
Table H.2 reports wall-clock runtimes (in seconds) for several NTK datasets, using the same models and configurations as in Table 2. It compares direct 64-bit computations performed on matrices stored in 16-, 32-, and 64-bit precisions to various approximations, including FLODANCE, SLQ, and others. All measurements were conducted on the same hardware under comparable conditions. FLODANCE consistently achieves sub-second runtimes even on subsamples of size $n=2500$, outperforming other approximations while avoiding the instability issues associated with SLQ.
Table H.2: Wall-clock runtimes (in seconds) for various log-determinant approximations ^$\hat{\ell}_{n}$, using the same models and configurations as in Table 2.
Quantity Model $\rightarrow$ Configuration ResNet9 ResNet9 ResNet18 MobileNet Dataset CIFAR-10 CIFAR-10 CIFAR-10 MNIST Subsample Size $n=1000$ $n=2500$ $n=1000$ $n=2500$ Runtime (seconds) Direct Computation (16-bit) $6.69$ $49.35$ $5.50$ $70.90$ Direct Computation (32-bit) $7.08$ $49.39$ $5.87$ 54.0554.05{54.05}54.05 Direct Computation (64-bit) $7.09$ $51.57$ $5.97$ $51.22$ SLQ $8.904$ $67.70$ $22.18$ $148.8$ Block Diagonal $0.003$ $0.009$ $0.003$ $0.008$ Pseudo NTK $0.015$ $0.077$ $0.014$ $0.082$ FLODANCE $n_{0}=1$,00 $n_{s}=50$ $0.008$ $0.008$ $0.008$ $0.009$ FLODANCE $n_{0}=1$,00 $n_{s}=100$ $0.017$ $0.018$ $0.016$ $0.017$ FLODANCE $n_{0}=300$, $n_{s}=500$ 0.3770.377{0.377}0.377 0.3500.350{0.350}0.350 0.3540.354{0.354}0.354 $0.376$
@@ Appendix I Implementation and Reproducibility Guide
We developed a Python package detkit333detkit is available for installation from PyPI (https://pypi.org/project/detkit), the documentation can be found at https://ameli.github.io/detkit, and the source code is available at https://github.com/ameli/detkit. that implements the MEMDET algorithm and can be used to reproduce the numerical results of this paper. A minimalistic usage of the detkit.memdet function is shown in LABEL:list:detkit1, where the user can specify various parameters: the maximum memory limit (max_mem), the structure of the matrix via the assume argument—set to gen for generic matrices (Algorithm 2), sym for symmetric matrices (Algorithm 3), and spd for symmetric positive-definite matrices (Algorithm 4)—whether the data is provided in full or in its lower/upper triangular form (triangle), the arithmetic precision used during computation (mixed_precision), the location of scratchpad space on disk (scratch_dir), and enabling parallel data transfer between memory and disk (parallel_io). The function in this example returns the log and sign of the determinant of the full-size matrix (ld, sign), along with the diagonal entries of matrix $\mathbf{D}$ (diag) and the array of permutation indices (perm) for the permutation matrix $\mathbf{P}$ from the LDL decomposition $\mathbf{P}^{\intercal}\mathbf{M}\mathbf{P}=\mathbf{L}\mathbf{D}\mathbf{L}^{\intercal}$.
@@
6M = zarr.open(’filename.zarr’, mode=’r’)
@@
The next example, shown in LABEL:list:detkit2, demonstrates the use of the FLODANCE method via the FitLogdet class in detkit. This method fits a scaling law to a small subset of log-determinants computed from submatrices and extrapolates to larger submatrix sizes using Algorithm 1. Specifically, we use the diag array from LABEL:list:detkit1 to compute the log-determinants $\ell_{k}=\sum_{i=1}^{kd}\log|D_{ii}|$, $k=n_{0},\dots,n_{s}$, for submatrices of size $kd\times kd$, where $d$ is the output dimension of the model (e.g., number of classes in CIFAR-10). These submatrices correspond to a permuted ordering of the original matrix during LDL decomposition, i.e., $\tilde{\mathbf{M}}\coloneqq\mathbf{P}^{\intercal}\mathbf{M}\mathbf{P}$. While the sampling of submatrices could, in principle, be performed in any order, the LDL decomposition conveniently provides the log-determinants of successive principal submatrices—formed by selecting the first $kd$ rows and columns—at no additional cost; an advantage we exploit in this approach. The resulting sequence $\ell_{k}$ is then fitted over the interval $k\in[n_{0},n_{s}]$, and the fitted FLODANCE model is used to predict log-determinants in the extrapolation range $[n_{s},n]$.
We have also concurrently developed a separate high-performance Python package, imate,444imate is available for installation from PyPI (https://pypi.org/project/imate), the documentation can be found at https://ameli.github.io/imate, and the source code is available at https://github.com/ameli/imate. which implements stochastic Lanczos quadrature (SLQ), a randomized method for approximating the log-determinant at scale. This package is implemented with a C++/CUDA backend and supports execution on both CPU and multiple GPUs. LABEL:list:imate demonstrates a minimalistic usage of the imate.logdet function.
@@
# Number of data (;#\$ n \$#;) and labels (;#\$ d \$#;)